# Channel Estimation for XL-MIMO Systems with Decentralized Baseband Processing: Integrating Local Reconstruction with Global Refinement

### v2

@@ Abstract
In this paper, we investigate the channel estimation problem for extremely large-scale multiple-input multiple-output (XL-MIMO) systems with a hybrid analog-digital architecture, implemented within a decentralized baseband processing (DBP) framework with a star topology. Existing centralized and fully decentralized channel estimation methods face limitations due to excessive computational complexity or degraded performance. To overcome these challenges, we propose a novel two-stage channel estimation scheme that integrates local sparse reconstruction with global fusion and refinement. Specifically, in the first stage, by exploiting the sparsity of channels in the angular-delay domain, the local reconstruction task is formulated as a sparse signal recovery problem. To solve it, we develop a graph neural networks-enhanced sparse Bayesian learning (SBL-GNNs) algorithm, which effectively captures dependencies among channel coefficients, significantly improving estimation accuracy. In the second stage, the local estimates from the local processing units (LPUs) are aligned into a global angular domain for fusion at the central processing unit (CPU). Based on the aggregated observations, the channel refinement is modeled as a Bayesian denoising problem. To efficiently solve it, we devise a variational message passing algorithm that incorporates a Markov chain-based hierarchical sparse prior, effectively leveraging both the sparsity and the correlations of the channels in the global angular-delay domain. Simulation results validate the effectiveness and superiority of the proposed SBL-GNNs algorithm over existing methods, demonstrating improved estimation performance and reduced computational complexity.
@@ I Introduction
To meet the increasing demand for data transmission, millimeter-wave (mmWave) bands have become essential to alleviate spectrum congestion in traditional microwave frequencies. However, high-frequency signals suffer from significant attenuation and are highly susceptible to blockages. To overcome these challenges, massive multiple-input multiple-output (MIMO) systems with hybrid beamforming architectures have been widely adopted to mitigate considerable path loss. Recently, the concept of extremely large-scale MIMO (XL-MIMO) has emerged as an evolution of the massive MIMO paradigm, attracting substantial attention from both academia and industry [1]. XL-MIMO systems are anticipated to achieve higher spectrum and energy efficiencies, expand network coverage, and enhance positioning accuracy by leveraging substantial spatial multiplexing and beamforming gains [2]. To fully realize these benefits, advanced signal processing techniques such as beamforming and equalization must be meticulously designed. Notably, obtaining accurate channel state information (CSI) is essential for realizing these performance gains, underscoring the critical role of precise channel knowledge in maximizing the potential of XL-MIMO technologies.
@@ I-A Existing Research
Recent research has extensively explored channel estimation algorithms tailored to various characteristics of XL-MIMO channels. For example, [3] addressed spherical wavefront effects by designing an angular-and-distance-dependent near-field codebook and proposed a polar-domain sparsity-based simultaneous orthogonal matching pursuit (SOMP) algorithm. This approach was further extended to hybrid near-field and far-field scenarios by [4, 5]. To mitigate the size of the polar-domain codebook, [6] introduced a distance-parameterized angular-domain representation, facilitating channel estimation through joint dictionary learning and sparse recovery methods.
@@
Considering the spatially non-stationary (SnS) property of XL-MIMO channels, [7] transformed SnS channels into a series of spatially stationary channels using a group time block code (GTBC)-based signal extraction scheme. Leveraging the antenna-domain sparsity of SnS channels, [8] characterized SnS channels with a subarray-wise Bernoulli-Gaussian prior and proposed a Bayesian inference-based scheme for joint user activity and channel estimation. Furthermore, to fully exploit antenna-and-delay-domain sparsity in SnS channels, [9] introduced a structured prior model based on a hidden Markov model (HMM) and proposed the turbo orthogonal approximate message passing (Turbo-OAMP) algorithm for low-complexity channel estimation. Furthermore, taking into account both antenna-domain-and-angular domain characteristics, [10] introduced a two-stage approach for detecting visibility regions (VRs) and estimating channels. However, the two stages are separate from each other, resulting in insufficient utilization of angular-domain information during the visibility region detection phase. To overcome the limitation, we further introduced a cascaded generalized approximate message passing (TL-GAMP) algorithm to jointly estimate VRs and channels in our subsequent work [11].
@@
Taking into account the near-field dual-wideband effect for wideband XL-MIMO systems, [12] proposed to model the angular-and-delay-domain channels with an independent and non-identically distributed Bernoulli-Gaussian prior, and the channel estimation problem is solved by the constrained Bethe free energy minimization framework. To further exploit the structure sparsity of angular-and-delay-domain channels, [13] introduced a 4-connected Markov random filed (MRF) based sparse prior model and proposed a unity approximate message passing to achieve the low-complexity estimation.
@@
However, the aforementioned algorithms are implemented centrally. As the number of antennas increases, such centralized architecture encounters bottlenecks in terms of high interconnection cost and computational complexity. To address these challenges, decentralized basedband processing (DBP) has emerged as a promising technique [14, 15, 16]. In the DBP architecture, the antennas are divided into several antenna clusters, and each cluster is processed by an independent and more affordable local processing unit (LPU). Notably, there are significant distinctions between the DBP architecture and cell-free systems. Specifically, in contrast to cell-free systems, where the access points (APs) are geographically distributed and potentially lack a centralized base station [17, 18], the DBP architecture retains a centralized antenna array. Furthermore, in the cell-free context, the APs have independent power constraints because they are geographically separated devices, each provisioned with its own power line. In contrast, the DBP network is a baseband signal processing layer architecture mounted on XL-MIMO antennas, where the segments share a common power supply[19].
@@ I-B Motivations and Contributions
  • To efficiently solve the sparse signal recovery problem in LPUs, we propose a graph neural networks-based sparse Bayesian learning (SBL-GNNs) algorithm. Specifically, the expectation step (E-step) of the traditional SBL algorithm is preserved, while the non-linear functions for the prior variance update in the maximization step (M-step) are replaced by GNNs. Based on the observations from the E-step, the prior variance update is formulated as a maximum a posteriori (MAP) estimation problem, modeled using a pairwise MRF. Leveraging the properties of this pairwise MRF, a GNN architecture is designed to efficiently capture dependencies among channel coefficients and perform MAP inference, enhancing both the accuracy and robustness of channel estimation.
@@
Organization: The rest of the paper is organized as follows. In Section II, we introduce the system model and problem description. Section III describes the proposed two-stage estimation scheme. In Section IV, we detail the proposed SBL-GNNs algorithm. Section V introduces the proposed central refinement algorithm for the CPU. Finally, simulation results are presented in Section VI, and conclusions are drawn in Section VII.
Notations: Lower-case letters, bold-face lower-case letters, and bold-face upper-case letters are used for scalars, vectors and matrices, respectively; The superscripts $\left(\cdot\right)^{\mathrm{T}}$ and $\left(\cdot\right)^{\mathrm{H}}$ stand for transpose and conjugate transpose, respectively; $\mathrm{blkdiag}\left(\mathbf{X}_{1},\mathbf{X}_{2},\dotso,\mathbf{X}_{N}\right)$ denotes a diagonal matrix with $\left\{\mathbf{X}_{1},\mathbf{X}_{2},\dotso,\mathbf{X}_{N}\right\}$ being its diagonal elements; $\odot$ denotes the Hadamard product; $\mathbb{C}^{M\times N}$ denotes a $M\times N$ complex matrix. In addition, a random variable $x\in\mathbb{C}$ drawn from the complex Gaussian distribution with mean $m$ and variance $v$ is characterized by the probability density function $\mathcal{CN}(x;m,v)={\exp\left\{-\{|x-m\right|^{2}}/{v}\}/{\pi v}$; a random variable $\gamma\in\mathbb{R}$ from Gamma distribution with mean $a/b$ and variance $a/b^{2}$ is characterized by $\mathcal{G}a(\gamma;a,b)\propto\gamma^{a-1}\exp(-\gamma b)$.
@@ II System Models and Problem Description
Consider a mmWave XL-MIMO system in which the base station (BS) equipped with an $N_{t}$-antenna extremely large-scale array (ELAA), arranged in a uniform linear configuration, serves $U$ single-antenna users. To enable efficient implementation, the BS adopts a hybrid analog-digital architecture based on a DBP framework with a star topology [24, 26], as shown in Fig. 1. In this configuration, the ELAA is divided into $M$ non-overlapping subarrays, each consisting of $N=N_{t}/M$ antennas. These subarrays are connected to LPUs through a phase shifter network, with each subarray associated with $N_{\mathrm{RF}}$ radio frequency (RF) chains. Then, the LPUs coordinate with a CPU through fronthaul links for baseband signal processing [^1] Due to the wired transmission of the fronthaul link, we assume that the different LPUs can achieve perfect synchronization..
Furthermore, to mitigate the effects of frequency-selective fading, orthogonal frequency-division multiplexing (OFDM) is employed [27, 28]. Without loss of generality, assume that $K$ subcarriers are uniformly selected from the available subcarriers to transmit pilot symbols. For notation conciseness later, we denote the following sets =$\mathcal{M}=\{1,2,\cdots,M\}$, $\mathcal{N}=\{1,2,\cdots,N\}$, $\mathcal{J}=\{1,2,\cdots,N_{t}\}$, and $\mathcal{K}=\{1,2,\cdots,K\}$.
@@ II-A Channel Model
$[\mathbf{a}(\tau_{g}^{(u)})]_{k}$ $=\mathrm{e}^{-\mathrm{j}2\pi{f_{k}}\tau_{g}^{(u)}},$ (2)
$[\mathbf{b}(\psi_{g}^{(u)},\varphi_{g}^{(u)})]_{n}$ $=\mathrm{e}^{-\mathrm{j}2\pi\left({\psi_{g}^{(u)}}(n-1)-{\varphi_{g}^{(u)}}(n-1)^{2}\right)},$ (3)
where $f_{k}=f_{c}+f_{s}(k-1)$ indicates the $k$-th subcarrier frequency with $f_{s}$ being the system bandwidth and $k\in\mathcal{K}$; $\tau_{g}^{(u)}=r_{g}^{(u)}/c$ denotes the path delay with $r_{g}^{(u)}$ indicating the distance between the reference antenna and user or scatters. In addition, $\psi_{g}^{(u)}\triangleq d\cos\vartheta_{g}^{(u)}/\lambda_{c}$ and $\varphi_{g}^{(u)}\triangleq{d^{2}\sin^{2}\vartheta_{g}^{(u)}}/{2r_{g}^{(u)}\lambda_{c}}$ with $\vartheta_{g}^{(u)}$ indicating the direction angle of the $g$-th path to the reference antenna; $\boldsymbol{\Theta}(\psi_{g}^{(u)},\varphi_{g}^{(u)})$ denotes the frequency-dependent phase shifts matrix, whose $(n,k)$-th element, $n\in\mathcal{N}$ and $k\in\mathcal{K}$, is given by
$\left[\boldsymbol{\Theta}(\psi_{g}^{(u)},\varphi_{g}^{(u)})\right]_{n,k}=\mathrm{e}^{-\mathrm{j}2\pi{f_{k}}\left((n-1)\frac{\psi_{g}^{(u)}}{f_{c}}-(n-1)^{2}\frac{\varphi_{g}^{(u)}}{f_{c}}\right)}.$ (4)
Consequently, the channel of user $u$ associated with subarray $m$ is represented by $\mathbf{H}_{m}^{(u)}=[\mathbf{h}_{m,1}^{(u)},\mathbf{h}_{m,2}^{(u)},\cdots,\mathbf{h}_{m,K}^{(u)}]\in\mathbb{C}^{N\times K}$, which consists of the rows of $\mathbf{H}^{(u)}$ from the $((m-1)N+1)$-th to the $mN$-th row.
@@ II-B Problem Description
This paper mainly addresses the uplink channel estimation task. Without loss of generality, we assume that the mutually orthogonal pilot sequences are used by different users. As a result, the channel estimation for each user is independent of the others. In this context, we focus on the channel estimation problem for an arbitrary user $u$. For notation simplification, we omit the superscript $u$ in the following. On the BS side, each subarray combines the received pilot signals using $N_{\mathrm{RF}}$ RF chains, each associated with a distinct beam. Assume the total number of receive beams as $Q$. As a result, we needs $P=Q/N_{\mathrm{RF}}$ time slots to traverse all receive beams given a fixed pilot symbol. Denote $\iota_{k,p}$ as the pilot symbol of the $k$-th subcarrier in the $p$-th time slot. Then, the received signal $\mathbf{y}_{m,k,p}\in\mathbb{C}^{N_{\mathrm{RF}}\times 1}$ of subarray $m$ at the $k$-th subcarrier in the $p$-th time slot is given by
@@
where $\mathbf{W}_{m,p}\in\mathbb{C}^{N_{\mathrm{RF}}\times N}$ denotes the combining matrix for subarray $m$ in the $p$-th time slot, with each element randomly selected from $\{-1/\sqrt{N},1/\sqrt{N}\}$; $\mathbf{h}_{m,k}\in\mathbb{C}^{N\times 1}$ and $\mathbf{n}_{m,k,p}\in\mathbb{C}^{N\times 1}$ indicate the channel and noise vectors, respectively, for subarray $m$ at the $k$-th subcarrier. Without loss the generality, we assume $\iota_{k,p}=1$ for all $k$ and $p$. By collecting the received pilot signals corresponding to all $P$ time slots, the total received signal $\mathbf{y}_{m,k}=[\mathbf{y}_{m,k,1}^{\mathrm{T}},\cdots,\mathbf{y}_{m,k,P}^{\mathrm{T}}]^{\mathrm{T}}\in\mathbb{C}^{Q\times 1}$ for subarray $m$ at the $k$-th subcarrier can be expressed as
@@
where $\mathbf{W}_{m}=\left[\mathbf{W}_{m,1}^{\mathrm{T}},\mathbf{W}_{m,2}^{\mathrm{T}},\cdots,\mathbf{W}_{m,P}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{C}^{Q\times N}$ and ${\mathbf{n}}_{m,k}=[\mathbf{n}_{m,k,1}^{\mathrm{T}}\mathbf{W}_{m,1}^{\mathrm{T}},\mathbf{n}_{m,k,2}^{\mathrm{T}}\mathbf{W}_{m,2}^{\mathrm{T}},\cdots,\mathbf{n}_{m,k,P}^{\mathrm{T}}\mathbf{W}_{m,P}^{\mathrm{T}}]^{\mathrm{T}}\in\mathbb{C}^{Q\times 1}$ are the collective combining matrix and the effective noise vector, respectively. Further, by concentrating the received signals from different subcarriers, the collective pilot signal $\mathbf{Y}_{m}=\left[\mathbf{y}_{m,1},\mathbf{y}_{m,2},\cdots,\mathbf{y}_{m,K}\right]\in\mathbb{C}^{Q\times K}$ at the $k$-th subcarrier is given by
@@
with $\mathbf{H}_{m}=\left[\mathbf{h}_{m,1},\mathbf{h}_{m,2},\cdots\mathbf{h}_{m,K}\right]\in\mathbb{C}^{N\times K}$ and $\mathbf{N}_{m}=\left[{\mathbf{n}}_{m,1},{\mathbf{n}}_{m,2},\cdots{\mathbf{n}}_{m,K}\right]\in\mathbb{C}^{Q\times K}$ indicating the collective channel matrix and noise matrix for subarray $m$, respectively.
@@
where $\mathbf{W}=\mathrm{blkdiag}\left(\mathbf{W}_{1},\mathbf{W}_{2},\cdots,\mathbf{W}_{M}\right)\in\mathbb{C}^{{QM\times N_{\mathrm{t}}}}$ and $\mathbf{N}=\left[\mathbf{N}_{1}^{\mathrm{T}},\mathbf{N}_{2}^{\mathrm{T}},\cdots,\mathbf{N}_{M}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{C}^{QM\times K}$ denote the overall received matrix and noise matrix, respectively. Utilizing the sparsity of $\mathbf{H}$ in the angular-delay domain [13], the channel estimation task can be formulated as sparse signal recovery problem, i.e.,

### v4

@@ Abstract
In this paper, we investigate the channel estimation problem for extremely large-scale multiple-input multiple-output (XL-MIMO) systems with a hybrid analog-digital architecture, implemented within a decentralized baseband processing (DBP) framework with a star topology. Existing centralized and fully decentralized channel estimation methods face limitations due to excessive computational complexity or degraded performance. To overcome these challenges, we propose a novel two-stage channel estimation scheme that integrates local sparse reconstruction with global fusion and refinement. Specifically, in the first stage, by exploiting the sparsity of channels in the angular-delay domain, the local reconstruction task is formulated as a sparse signal recovery problem. To solve it, we develop a graph neural networks-enhanced sparse Bayesian learning (SBL-GNNs) algorithm, which effectively captures dependencies among channel coefficients, significantly improving estimation accuracy. In the second stage, the local estimates from the local processing units (LPUs) are aligned into a global angular domain for fusion at the central processing unit (CPU). Based on the aggregated observations, the channel refinement is modeled as a Bayesian denoising problem. To efficiently solve it, we devise a variational message passing algorithm that incorporates a Markov chain-based hierarchical sparse prior, effectively leveraging both the sparsity and the correlations of the channels in the global angular-delay domain. Simulation results show the effectiveness and superiority of the proposed SBL-GNNs algorithm over existing methods, demonstrating improved estimation performance and reduced computational complexity.
@@ I Introduction
To meet the increasing demand for data transmission, millimeter-wave (mmWave) bands have become essential to alleviate spectrum congestion in traditional microwave frequencies. However, high-frequency signals suffer from significant attenuation and are highly susceptible to blockages. To overcome these challenges, massive multiple-input multiple-output (MIMO) systems with hybrid beamforming architectures have been widely adopted to mitigate considerable path loss. Recently, the concept of extremely large-scale MIMO (XL-MIMO) has emerged as an evolution of the massive MIMO paradigm, attracting substantial attention from both academia and industry [1]. XL-MIMO systems are anticipated to achieve higher spectrum and energy efficiencies, expand network coverage, and enhance positioning accuracy by leveraging substantial spatial multiplexing and beamforming gains [2, 3, 4]. To fully realize these benefits, advanced signal processing techniques such as beamforming and equalization must be meticulously designed. Notably, obtaining accurate channel state information (CSI) is essential for realizing these performance gains, underscoring the critical role of precise channel knowledge in maximizing the potential of XL-MIMO technologies [5, 6].
@@ I-A Existing Research
Recent research has extensively explored channel estimation algorithms tailored to various characteristics of XL-MIMO channels. For example, [7] addressed spherical wavefront effects by designing an angular-and-distance-dependent near-field codebook and proposed a polar-domain sparsity-based simultaneous orthogonal matching pursuit (SOMP) algorithm. This approach was further extended to hybrid near- and far-field scenarios by [8, 9]. To further mitigate the size of the polar-domain codebook, [10] introduced a distance-parameterized angular-domain representation, facilitating channel estimation through joint dictionary learning and sparse recovery methods.
@@
Considering the spatially non-stationary (SnS) property of XL-MIMO channels, [11] transformed SnS channels into a series of spatially stationary channels using a group time block code (GTBC)-based signal extraction scheme. Leveraging the antenna-domain sparsity of SnS channels, [12] characterized SnS channels with a subarray-wise Bernoulli-Gaussian prior and proposed a Bayesian inference-based scheme for joint user activity and channel estimation. Furthermore, to fully exploit the sparsity of SnS channels in the antenna-delay domain, [13] introduced a structured prior model based on a hidden Markov model (HMM) and proposed the turbo orthogonal approximate message passing (Turbo-OAMP) algorithm for low-complexity channel estimation. Furthermore, taking into account both antenna-domain-and-angular domain characteristics, [14] introduced a two-stage approach for detecting visibility regions (VRs) and estimating channels. However, the two stages are performed separately, leading to insufficient utilization of angular-domain information during the VRs detection phase. To overcome the limitation, we further introduced a cascaded generalized approximate message passing (TL-GAMP) algorithm to jointly estimate VRs and channels in our subsequent work [15].
@@
Taking into account the near-field dual-wideband effect for wideband XL-MIMO systems, [16] proposed to model the angular-delay domain channels with an independent and non-identically distributed Bernoulli-Gaussian prior, and the channel estimation problem is solved by the constrained Bethe free energy minimization framework. To further exploit the structure sparsity of angular-delay domain channels, [17] introduced a 4-connected Markov random filed (MRF) based sparse prior model and proposed a unity approximate message passing to achieve the low-complexity estimation.
@@
However, the aforementioned algorithms are implemented centrally. As the number of antennas increases, such centralized architecture encounters bottlenecks in terms of high interconnection cost and computational complexity. To address these challenges, decentralized baseband processing (DBP) has emerged as a promising technique [18, 19, 20]. In the DBP architecture, the antennas are divided into several antenna clusters, and each cluster is processed by an independent and more affordable local processing unit (LPU). Notably, there are significant distinctions between the DBP architecture and cell-free systems. Specifically, in contrast to cell-free systems, where the access points (APs) are geographically distributed and potentially lack a centralized base station [21, 22, 23], the DBP architecture retains a centralized antenna array. Furthermore, in the cell-free context, the APs have independent power constraints because they are geographically separated devices, each provisioned with its own power line. In contrast, the DBP network is a baseband signal processing layer architecture mounted on XL-MIMO antennas, where the segments share a common power supply[24].
@@ I-B Motivations and Contributions
  • To efficiently solve the sparse signal recovery problem in LPUs, we propose a graph neural networks-based sparse Bayesian learning (SBL-GNNs) algorithm. Specifically, the expectation step (E-step) of the traditional SBL algorithm is preserved, while the non-linear functions for the prior variance update in the maximization step (M-step) are replaced by GNNs. Based on the observations from the E-step, the prior variance update is formulated as a maximum a posteriori (MAP) estimation problem, modeled using a pairwise MRF representation. Leveraging the properties of this pairwise MRF, a GNN architecture is designed to efficiently capture dependencies among channel coefficients and perform MAP inference, enhancing both the accuracy and robustness of channel estimation.
@@
Organization: The rest of the paper is organized as follows. In Section II, we introduce the system model and problem description. Section III describes the proposed two-stage estimation scheme. In Section IV, we detail the proposed SBL-GNNs algorithm. Section V introduces the proposed centralized refinement algorithm for the CPU. Finally, simulation results are presented in Section VI, and conclusions are drawn in Section VII.
Notations: Lower-case letters, bold-face lower-case letters, and bold-face upper-case letters are used for scalars, vectors and matrices, respectively; The superscripts $\left(\cdot\right)^{\mathrm{T}}$ and $\left(\cdot\right)^{\mathrm{H}}$ stand for transpose and conjugate transpose, respectively; $\mathrm{blkdiag}\left(\mathbf{X}_{1},\mathbf{X}_{2},\dotso,\mathbf{X}_{N}\right)$ denotes a diagonal matrix with $\left\{\mathbf{X}_{1},\mathbf{X}_{2},\dotso,\mathbf{X}_{N}\right\}$ being its diagonal elements; $\otimes$ denotes the Kronecker product; $\mathrm{vec}(\cdot)$ denotes the vectorization operations, which reorganizes the matrix into a column vector. $\mathbb{C}^{M\times N}$ denotes a $M\times N$ complex matrix. In addition, a random variable $x\in\mathbb{C}$ drawn from the complex Gaussian distribution with mean $m$ and variance $v$ is characterized by the probability density function $\mathcal{CN}(x;m,v)={\exp\left\{-\{|x-m\right|^{2}}/{v}\}/{\pi v}$; a random variable $\gamma\in\mathbb{R}$ from Gamma distribution with mean $a/b$ and variance $a/b^{2}$ is characterized by $\mathcal{G}a(\gamma;a,b)\propto\gamma^{a-1}\exp(-\gamma b)$.
@@ II System Models and Problem Description
Consider a mmWave XL-MIMO system in which the base station (BS) equipped with an $N_{t}$-antenna extremely large-scale array (ELAA), arranged in a uniform linear configuration, serves $U$ single-antenna users. Additionally, to enable efficient implementation, the BS adopts a hybrid analog-digital architecture based on a DBP framework with a star topology [30, 31], as shown in Fig. 1. In this configuration, the ELAA is divided into $M$ non-overlapping subarrays, each consisting of $N=N_{{t}}/M$ antennas. Without loss of generality, we assume the antenna spacing is $d=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$ with $c$ and $f_{c}$ indicating the speed of light and central carrier frequency, respectively. These subarrays are connected to LPUs through a phase shifter network, with each subarray associated with $N_{\mathrm{RF}}$ radio frequency (RF) chains. Then, the LPUs coordinate with a CPU through fronthaul links for baseband signal processing [^1] Due to the wired transmission of the fronthaul link, we assume that the different LPUs can achieve perfect synchronization..
Furthermore, to mitigate the effects of frequency-selective fading, orthogonal frequency-division multiplexing (OFDM) is employed. Without loss of generality, assume that $K$ subcarriers are uniformly selected from the available subcarriers to transmit pilot symbols. For notation conciseness later, we denote the following sets =$\mathcal{M}=\{1,2,\cdots,M\}$, $\mathcal{N}=\{1,2,\cdots,N\}$, $\mathcal{J}=\{1,2,\cdots,N_{t}\}$, $\mathcal{K}=\{1,2,\cdots,K\}$, and $\mathcal{V}=\{1,2,\cdots,NK\}$.
@@ II-A Channel Model
$[\mathbf{a}(\tau_{g}^{(u)})]_{k}$ $=\mathrm{e}^{-\mathrm{j}2\pi{f_{k}}\tau_{g}^{(u)}},k\in\mathcal{K}$ (2)
$[\mathbf{b}(\psi_{g}^{(u)},\varphi_{g}^{(u)})]_{j}$ $=\mathrm{e}^{-\mathrm{j}2\pi\left({\psi_{g}^{(u)}}(j-1)-{\varphi_{g}^{(u)}}(j-1)^{2}\right)},j\in\mathcal{J}$ (3)
where $f_{k}=f_{c}+f_{s}(k-1)$ indicates the $k$-th subcarrier frequency with $f_{s}$ being the system bandwidth; $\tau_{g}^{(u)}=r_{g}^{(u)}/c$ denotes the path delay with $r_{g}^{(u)}$ indicating the distance between the reference antenna and user or scatters. In addition, $\psi_{g}^{(u)}\triangleq d\cos\vartheta_{g}^{(u)}/\lambda_{c}$ and $\varphi_{g}^{(u)}\triangleq{d^{2}\sin^{2}\vartheta_{g}^{(u)}}/{2r_{g}^{(u)}\lambda_{c}}$ with $\vartheta_{g}^{(u)}$ indicating the direction angle of the $g$-th path to the reference antenna. From (3), the array steering vectors $\mathbf{b}(\psi_{g},\varphi_{g})$ are distance-dependent and incorporate quadratic terms with respect to the antenna index $n$, accounting for variations in path delay and direction angle across antenna elements, thereby reflecting the impact of near-field propagation. $\boldsymbol{\Theta}(\psi_{g}^{(u)},\varphi_{g}^{(u)})$ denotes the frequency-dependent phase shifts matrix, whose $(j,k)$-th element, $j\in\mathcal{J}$ and $k\in\mathcal{K}$, is given by
$\left[\boldsymbol{\Theta}(\psi_{g}^{(u)},\varphi_{g}^{(u)})\right]_{j,k}=\mathrm{e}^{-\mathrm{j}2\pi{f_{k}}\left((j-1)\frac{\psi_{g}^{(u)}}{f_{c}}-(j-1)^{2}\frac{\varphi_{g}^{(u)}}{f_{c}}\right)}.$ (4)
Consequently, the channel of user $u$ associated with subarray $m$ is represented by $\mathbf{H}_{m}^{(u)}=[\mathbf{h}_{m,1}^{(u)},\mathbf{h}_{m,2}^{(u)},\cdots,\mathbf{h}_{m,K}^{(u)}]\in\mathbb{C}^{N\times K}$, which consists of the rows of $\mathbf{H}^{(u)}$ from the $((m-1)N+1)$-th to the $mN$-th row with $m\in\mathcal{M}$.
@@ II-B Problem Description
This paper mainly addresses the uplink channel estimation task. Without loss of generality, we assume that the mutually orthogonal pilot sequences are used by different users. As a result, the channel estimation for each user is independent of the others. In this context, we focus on the channel estimation problem for an arbitrary user $u$. For notation simplification, we omit the superscript $u$ in the following. At the BS side, each subarray combines the received pilot signals using $N_{\mathrm{RF}}$ RF chains, each associated with a distinct beam. Assume the total number of training beams as $Q$. As a result, we needs $P=Q/N_{\mathrm{RF}}$ time slots to traverse all receive beams given a fixed pilot symbol. Denote $\iota_{k,p}$ as the pilot symbol of the $k$-th subcarrier in the $p$-th time slot. Then, the received signal $\mathbf{y}_{m,k,p}\in\mathbb{C}^{N_{\mathrm{RF}}\times 1}$ of subarray $m$ at the $k$-th subcarrier in the $p$-th time slot is given by
@@
where $\mathbf{W}_{m,p}\in\mathbb{C}^{N_{\mathrm{RF}}\times N}$ denotes the combining matrix for subarray $m$ in the $p$-th time slot, $m\in\mathcal{M}$ and $p=1,2,\cdots,P$, with each element randomly selected from $\{-1/\sqrt{N},1/\sqrt{N}\}$; $\mathbf{h}_{m,k}\in\mathbb{C}^{N\times 1}$ and $\mathbf{n}_{m,k,p}\in\mathbb{C}^{N\times 1}$ indicate the channel and noise vectors, respectively, for subarray $m$ at the $k$-th subcarrier with $m\in\mathcal{M}$ and $k\in\mathcal{K}$. Additionally, $\mathbf{n}_{m,k,p}$ is the additional white Gaussian (AWGN) noise vector, where each element obeys circularly symmetric complex Gaussian distribution with variance $\zeta_{m}^{-1}$ with $\zeta_{m}$ indicating the variance precision parameters. Without loss the generality, we assume $\iota_{k,p}=1$ for all $k$ and $p$. By collecting the received pilot signals corresponding to all $P$ time slots, the total received signal $\mathbf{y}_{m,k}=[\mathbf{y}_{m,k,1}^{\mathrm{T}},\cdots,\mathbf{y}_{m,k,P}^{\mathrm{T}}]^{\mathrm{T}}\in\mathbb{C}^{Q\times 1}$ for subarray $m$ at the $k$-th subcarrier can be expressed as
@@
where $\mathbf{W}_{m}=\left[\mathbf{W}_{m,1}^{\mathrm{T}},\mathbf{W}_{m,2}^{\mathrm{T}},\cdots,\mathbf{W}_{m,P}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{C}^{Q\times N}$ is the collective combining matrix. ${\mathbf{n}}_{m,k}=[\mathbf{n}_{m,k,1}^{\mathrm{T}}\mathbf{W}_{m,1}^{\mathrm{T}},\mathbf{n}_{m,k,2}^{\mathrm{T}}\mathbf{W}_{m,2}^{\mathrm{T}},\cdots,\mathbf{n}_{m,k,P}^{\mathrm{T}}\mathbf{W}_{m,P}^{\mathrm{T}}]^{\mathrm{T}}\in\mathbb{C}^{Q\times 1}$ is the effective noise vector with the covariance matrix $\mathrm{blkdiag}(\zeta_{m}^{-1}\mathbf{W}_{m,1}\mathbf{W}_{m,1}^{\mathrm{H}},\cdots,\zeta_{m}^{-1}\mathbf{W}_{m,P}\mathbf{W}_{m,P}^{\mathrm{H}})$. It follows that the incorporation of the analog combiner matrix $\mathbf{W}_{m,p}$ can lead to a non-identity covariance matrix $\mathbf{R}$, indicating that $\mathbf{n}_{m,k}$ may exhibit colored characteristics. Notably, if $\mathbf{W}_{m,p}$ is a semi-unitary matrix for any $m$ and $p$, i.e., $\mathbf{W}_{m,p}\mathbf{W}_{m,p}^{\mathrm{H}}=\mathbf{I}_{N_{\mathrm{RF}}}$, the covariance matrix $\mathbf{R}$ simplifies to $\zeta_{m}^{-1}\mathbf{I}_{Q}$. In this case, $\mathbf{n}_{m,k}$ remains strictly white noise and follows $\mathbf{n}_{m,k}\sim\mathcal{CN}(\mathbf{0},\zeta_{m}^{-1}\mathbf{I}_{Q})$. Consequently, for simplicity, this manuscript assumes $\mathbf{W}_{m,p}$ to be semi-unitary for all $m$ and $p$. In practical implementations, $\mathbf{W}_{m,p}$ can be chosen as a partial DFT or Hadamard matrix. [^2] Notably, this assumption does not affect the subsequent algorithm design. The proposed estimation scheme is also applicable for a non-orthogonal analog combiner matrix, requiring only a pre-whitening to (11) [15]. Further, by concentrating the received signals from different subcarriers, the collective pilot signal $\mathbf{Y}_{m}=\left[\mathbf{y}_{m,1},\mathbf{y}_{m,2},\cdots,\mathbf{y}_{m,K}\right]\in\mathbb{C}^{Q\times K}$ at the $k$-th subcarrier is given by
@@
where $\mathbf{H}_{m}=\left[\mathbf{h}_{m,1},\mathbf{h}_{m,2},\cdots,\mathbf{h}_{m,K}\right]\in\mathbb{C}^{N\times K}$ and $\mathbf{N}_{m}=\left[{\mathbf{n}}_{m,1},{\mathbf{n}}_{m,2},\cdots,{\mathbf{n}}_{m,K}\right]\in\mathbb{C}^{Q\times K}$ represent the collective channel matrix and noise matrix for subarray $m$, respectively. In addition, each column of ${\mathbf{N}}_{m}$ is a white noise vector with identical covariance matrix $\zeta_{m}^{-1}\mathbf{I}_{Q}$.
@@
where $\mathbf{W}=\mathrm{blkdiag}\left(\mathbf{W}_{1},\mathbf{W}_{2},\cdots,\mathbf{W}_{M}\right)\in\mathbb{C}^{{QM\times N_{{t}}}}$ and $\mathbf{N}=\left[\mathbf{N}_{1}^{\mathrm{T}},\mathbf{N}_{2}^{\mathrm{T}},\cdots,\mathbf{N}_{M}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{C}^{QM\times K}$ denote the overall received matrix and noise matrix, respectively. Utilizing the sparsity of $\mathbf{H}$ in the angular-delay domain [17], the channel estimation task can be formulated as sparse signal recovery problem, i.e.,